# Continuity and equivariant dimension

### v2

@@ 1 Introduction
In the general noncommutative setting, key elements of the above discussion translate into conditions on freeness or equivariant dimension. Precursors to the study of equivariant dimension appeared in [56, §3], which motivated some later problems in noncommutative Borsuk-Ulam theory. Taghavi’s questions primarily concerned counting the number of odd self-adjoint elements required to produce an invertible square-sum. It should be noted that this is not the only interpretation of noncommutative Borsuk-Ulam theory; references [61, 3, 39, 41, 10, 11] consider an approach focused on (non-)existence of equivariant morphisms. The full formulation of equivariant dimension, including a much more general context, appeared in [19]. The local-triviality dimension of an action is defined in [19, Definition 3.1], and it is a noncommutative analogue of the Schwarz genus. Moreover, it is related to the Rokhlin dimension of [23, Definition 1.1], as explored in [19, §5.2]. We recall the definitions of the local-triviality dimension in Section 3.
In this manuscript, we will typically focus on actions of ${\mathbb{S}}^{1}$ or its finite cyclic subgroups, for which equivalent versions of the dimensions we study are found in [12, §3]. Consider, for example, a unital $C^{*}$-algebra $A$ on which ${\mathbb{Z}}/2$ acts. The action decomposes $A$ into eigenspaces for eigenvalues ±$\pm 1$; elements in the $-1$ eigenspace are called odd and elements in the $1$ eigenspace are called even. The local-triviality (LT for short) dimension of this action is
@@
where we note that the value $\infty$ is possible. With this notation, the Borsuk-Ulam Theorem is equivalent to the claim that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(C({\mathbb{S}}^{n}))=n$, where ${\mathbb{Z}}/2$ acts on $C({\mathbb{S}}^{n})$ via the antipodal action. Because $C({\mathbb{S}}^{n})$ is commutative, there are adjustments to this dimension that make no difference here, but vastly change the computation for noncommutative $C^{*}$-algebras. For example, one could relax the condition on the sum-square of the $a_{i}$ to be merely invertible instead of $1$. Alternatively, one could strengthen the requirements by insisting that the $a_{i}$ are normal commuting elements of $A$. Of particular note is that finiteness of the local-triviality dimension implies freeness [19, Theorem 3.8], though this is not always equivalent.
Here, we will study how the local-triviality dimension and its variants behave with respect to deformation and fields of $C^{*}$-algebras, with a focus on noncommutative spheres and tori. We are interested in circumstances where the local-triviality dimensions are continuous, or not continuous, with respect to a deformation parameter. By construction of examples, we find that the most one can generally hope for is semicontinuity of the dimension, as explored in Section 3. Further, we extend work of [12] by finding free actions whose weak local-triviality dimension is infinite. In Section 4, we explore the $\theta$-deformed spheres and tori in more detail, with a focus on the rational case. In particular, viewing a rational noncommutative torus as a space of sections of a finite-dimensional vector bundle yields information about the strong local-triviality dimension. These results strengthen the noncommutative Borsuk-Ulam theory of [39] when the domain sphere is commutative.
@@ 2 Preliminaries
For what follows, $\mathrm{Prim}(A)$ denotes the primitive spectrum [43, §4.1.2] of a $C^{*}$-algebra $A$: the space of primitive ideals, equipped with the Jacobson (or hull-kernel) topology [43, §4.1.4]. Further, $Z(A)$ denotes the center of $A$, and $M(A)$ denotes the multiplier algebra of $A$. Following [46, §3], say, we write
@@
The Natsume-Olsen spheres, which we will also call $\theta$-spheres for brevity, have $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\cong{\mathbb{Z}}$ [35], with a specific generator that is inspired by the commutative case.
@@
Recall (e.g. [35, p.1094]) that for the classical 3-sphere, the group $K_{1}(C({\mathbb{S}}^{3}))$ is generated by the element
@@
More generally, [35, §5] shows that the cyclic group $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\cong{\mathbb{Z}}$ admits a generator of matrix dimension $2^{n-1}$, which we will denote $Z_{\theta}$, such that each entry is a *-monomial. For each of the ${\mathbb{Z}}/k$ rotation actions defined above, the eigenspace of each entry is such that there exist unitary matrices $A,B\in M_{2^{n-1}}({\mathbb{C}})$, independent of $\theta$, such that $R(Z_{\theta})=AZ_{\theta}B$ (where $R$ is applied entrywise) and $A^{k}=B^{k}=I$. The antipodal action ($k=2$) is particularly simple; in this case each entry of $Z_{\theta}$ is odd, so $R(Z_{\theta})=-Z_{\theta}$. See [39, §3] for a slightly modified construction that chooses the monomial coefficients continuously.
The Borsuk-Ulam Theorem extends to the Natsume-Olsen spheres by [39, Corollary 4.12]; if $\phi:C({\mathbb{S}}^{2n-1}_{\theta})\to C({\mathbb{S}}^{2n-1}_{\omega})$ is ${\mathbb{Z}}/k$-equivariant, then the $K_{1}$ degree $m$ of $\phi$ is congruent to $1$ modulo $k$. In particular, the induced map $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\to K_{1}(C({\mathbb{S}}^{2n-1}_{\omega}))$ is nontrivial. However, the local triviality dimensions of these actions are not all identical. In particular, [10, Proposition 3.21] shows that the $\theta$-sphere $C({\mathbb{S}}^{2n-1}_{\theta})$ such that each generator anticommutes has local-triviality dimension $1$. Since the local-triviality dimension is integer-valued and the commutative sphere has dimension $12n-$ 1, this immediately shows that the local-triviality dimension fails to be continuous in the parameter $\theta$.
The $\theta$-spheres are related to a similar family of $C^{*}$-algebras called quantum (or noncommutative) tori ([20, §12.2] or [48, §1]). If $\theta$ is an antisymmetric $n\times n$ real matrix, the $C^{*}$-algebra $A^{n}_{\theta}$, or $C(\mathbb{T}^{n}_{\theta})$, is defined by generators $u_{1},\ldots,u_{n}$ and relations
@@
Since the set of $2$-dimensional antisymmetric matrices is one-dimensional, we abuse notation somewhat when $n=2$, and simply refer to $\theta=\theta_{12}\in\mathbb{R}$. The noncommutative tori and noncommutative spheres are related by polar decomposition [35, Theorem 2.5], which we recall in Section 4. In particular, the action $R$ discussed on $\theta$-spheres factors through a related action on the quantum tori:
@@
All $\theta$-spheres and noncommutative tori are known to be Rieffel deformations in the sense of [50, §10.2]. While it is not clear if the $\theta$-spheres or noncommutative tori form a continuous field over the set of antisymmetric matrices, [50, Theorem 8.13] implies that the field is separately continuous in each entry. In particular, the distinction between those two cases does not matter for $3$-spheres or $2$-tori. As we will discuss continuity of fields in more detail, we recall some basic definitions.
Recall [25, Definition 1.5] that for a locally compact Hausdorff space $X$, a $C_{0}(X)$-algebra is a $C^{*}$-algebra $A$ equipped with a non-degenerate morphism $C_{0}(X)\to Z(M(A))$. For us, $X$ will typically be compact and $A$ will be unital, so there will be little need to worry about units. We will also revert to the notation $C(X)=C_{0}(X)$ when $X$ is compact.
@@
For a group ${\mathbb{G}}$ and a locally compact Hausdorff space $X$, a ${\mathbb{G}}$-$C_{0}(X)$-algebra is a $C_{0}(X)$-algebra equipped with an action of ${\mathbb{G}}$ by automorphisms that fix the structure morphism $C_{0}(X)\to Z(M(A))$. $\blacklozenge$
$C(X)$-algebras are one way to formalize the notion of a field of $C^{*}$-algebras [47, Definition 1.1 and Propositions 1.2 and 1.3], while [37, Theorem 2.3] equates the notion to that of an upper semicontinuous $C^{*}$-algebra bundle, generalizing [18, Definition VII.8.2].
@@
A $C(X)$-algebra $A$ is called a continuous field if for each $a\in A$, the mapping
@@ 3 Local-Triviality Dimension and Fields
If $(C({\mathbb{G}}),\Delta)$ is a compact quantum group with a coaction on a unital $C^{*}$-algebra $A$, then [19, Definition 3.1] defines the weak and plain local-triviality dimensions as follows. Here, we let $t$ denote the identity function in $C_{0}((0,1])$. We also note that since $C({\mathbb{G}})$ denotes a compact quantum group, it need not be commutative as a $C^{*}$-algebra, despite the notation used.
@@
The local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{LT}}(A)$ is the infinimum of the set of $n\in\mathbb{N}$ such that there exist ${\mathbb{G}}$-equivariant *-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)=1$. $\blacklozenge$
@@
The weak local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{WLT}}(A)$ is the infinimum of the set of $n\in\mathbb{N}$ such that there exist ${\mathbb{G}}$-equivariant *-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)$ invertible. $\blacklozenge$
The strong variant of the local-triviality dimension, found in [19, Definition 3.20], is defined in reference to the noncommutative join operation of [14, Definition 4.1] or [3, Definition 1.4]. Here $E_{n}^{\Delta}{\mathbb{G}}$ denotes the join of $n+1$ copies of ${\mathbb{G}}$.
@@
A coaction $\delta:A\to A\otimes_{\text{min}}C({\mathbb{G}})$ is free if
@@
Multiple characterizations of freeness are also found in [4, Theorem 0.4]. Moreover, finiteness of any one of the local-triviality dimensions implies freeness of the coaction by [19, Theorem 3.8]. We will primarily be interested in the case that $A$ is noncommutative, but ${\mathbb{G}}$ is an ordinary group (in particular, a compact abelian group). In this case, there is an equivalent condition to freeness found in [49, Definition 1.6] or [44, Definition 7.1.4]. If ${\mathbb{G}}$ is compact abelian and $\alpha$ is an action of ${\mathbb{G}}$ on $A$, then for each character $\gamma\in\widehat{{\mathbb{G}}}$, there is an associated eigenspace (or spectral subspace)
@@
This “asterisk” shape is star-convex, with center 00, generated by the $k$th roots of unity.
@@
Since $\operatorname{Star}_{2}=[-1,1]$, the elements $a_{i}$ used to compute ${\mathbb{Z}}/2$-dimensions are self-adjoints. As such, dimensions of ${\mathbb{Z}}/2$-actions are typically much easier to compute than other cases. Also, as the set of invertible elements of $A$ is open, but $\{1\}$ is closed, the propositions suggest that weak and plan local-triviality dimensions may behave differently for fields of $C^{*}$-algebras. At the very least, for ${\mathbb{G}}$-$C(X)$-algebras, we have the following immediate estimate on the global dimensions in terms of their respective fiber-wise counterparts.
@@
For a compact group ${\mathbb{G}}$ and a ${\mathbb{G}}$-$C(X)$-algebra $A$, we have
@@
The $({\mathbb{Z}}/2)$-action, on the other hand, is “twisted” conformant to the topology of $X$: the action on the fiber $A_{x}\cong M_{2}$ for a line
@@
so the $1$-eigenspace of $v_{x}$ will coincide with neither of the two eigenspaces of $u_{x}$ (with respective eigenvalues ±$\pm 1$). Identifying
@@
said eigenspaces of $u_{x}$ are nothing but $x$ and its antipodal point $-x\in{\mathbb{S}}^{2}$. The map
@@
is thus a self-map of the sphere that has no fixed points and sends no point to its antipode, and no such map exists by [58, Corollary 1.24]. $\blacklozenge$
@@
Equip ${\mathbb{P}}^{1}$ with its complex analytic structure or, equivalently for the purpose of manipulating holomorphic bundles, its complex projective algebraic structure (see [53, §12]). The line bundle
$X={\mathbb{P}}^{1}={\mathbb{P}}^{1}\ni x\xmapsto{\quad}\text{$1$-eigenspace of}u_{x}$
is the universal subbundle [16, §3.2.3] on ${\mathbb{P}}^{1}$ and hence what algebraic geometers denote by ${\mathcal{O}}(-1)$ [21, Definition preceding Proposition 5.12]. Its orthogonal complement
@@
in the trivial rank-2 bundle on ${\mathbb{P}}^{1}$ is then ${\mathcal{O}}(1)$, as the degrees [21, Exercise II.6.12] of the two bundles must add up to 00. A choice of non-zero self-adjoint odd $v_{x}$ for $x$ ranging over some open $U\subseteq{\mathbb{P}}^{1}$ means precisely an identification ${\mathcal{O}}(-1)\cong{\mathcal{O}}(1)$ over $U$, or a trivialization of
@@
Such trivializations exist over any proper open subset of ${\mathbb{P}}^{1}$ (e.g. the complement of a point), per the standard classification [38, Theorem 2.1.1] of line bundles on ${\mathbb{P}}^{1}$. It follows that there are trivializations
@@
A salient example of a $C(X)$-algebra is that of sections of an $n\times n$ matrix-algebra bundle ${\mathcal{M}}$ on $X$ in the sense of [24, Definition 18.1.3]: a vector bundle ${\mathcal{M}}\to X$ with local trivializations ${\mathcal{M}}|_{U}\cong U\times M_{n}$ respecting the $C^{*}$-algebra structures along the fibers. There will be no harm in allowing $n$ to vary, as $X$ simply decomposes into clopen subspaces over which $n$ is constant.
Having to work with such trivializations extensively, we need some language to express the extent to which they exist. To that end, the following notions specialize the (numerical) ${\mathbb{G}}$-index ([62, Definition 3.13], [30, Definition 6.2.3]) of a ${\mathbb{G}}$-action on a space $X$, which is typically compact. That invariant can be recast via [13, Theorem 3.4] (where the term is co-index instead) as the smallest $n$ for which $X/{\mathbb{G}}$ admits an open cover by $U_{0},\ldots,U_{n}$, above which the action induces trivial principal ${\mathbb{G}}$-bundles.
@@
    Parts 1 and 2 of Definition 3.9 are indeed instances of numerical ${\mathbb{G}}$-indices (at least along the clopen subsets of $X$ where the ranks of the vector/matrix bundles involved are constant), with ${\mathbb{G}}$ either a unitary or projective unitary group. As explained in [24, Assertion 18.3.4], rank-$n$ vector bundles and $n\times n$ matrix bundles “are” nothing but principal $U(n)$- and $PU(n)$-bundles, respectively.
@@
Because $\operatorname{\mathrm{qt}}$ is assumed finite, each ${\mathcal{E}}_{i,j}$ has nonzero vector space dimension at each $x\in X$. If $i$ is fixed and $x\in U_{i}$ is fixed, consider the smallest such eigenspace $W_{0}:=({\mathcal{E}}_{i,j_{0}})_{x}$. It is possible to define a partial isometry on $M_{n_{x}}$ that maps $W_{0}$ isometrically onto a subspace $W_{1}\subseteq({\mathcal{E}}_{i,j_{0}+1})_{x}$ (which is guaranteed to be orthogonal to $W_{0}$ because the eigenspaces are orthogonal), maps $W_{1}$ isometrically onto some $W_{2}\subseteq({\mathcal{E}}_{i,j_{0}+2})_{x}$, and so on, eventually returning to $W_{k}=W_{0}=({\mathcal{E}}_{i,j_{0}})_{x}$. Further, map every element of $\left(\bigoplus_{m=0}^{k-1}W_{m}\right)^{\perp}$ to 00. This procedure produces a partial isometry $\alpha$ in the generating spectral subspace of the action of ${\mathbb{Z}}/k$ on this fiber. In fact, $\alpha$ is normal, as the initial space and final space are both equal to the direct sum $\bigoplus_{m=0}^{k-1}W_{m}$, and it also follows from the construction that $\alpha^{k}$ is equal to the projection onto that same direct sum. The subspaces $W_{m}$ for $m>0$ may be proper subspaces of the corresponding eigenspaces, but because $\operatorname{\mathrm{qt}}$ is defined as the (largest) quotient spread, we need at most $\operatorname{\mathrm{qt}}$ such operators $\alpha_{1},\ldots,\alpha_{\operatorname{\mathrm{qt}}}$ to guarantee full coverage of each eigenspace. That is, $\sum\limits_{s=1}^{\operatorname{\mathrm{qt}}}\alpha_{s}^{*}\alpha_{s}$ is an invertible element of $M_{n_{x}}$.

### v4

@@ 1 Introduction
In the general noncommutative setting, key elements of the above discussion translate into conditions on freeness or equivariant dimension. Precursors to the study of equivariant dimension appeared in [57, §3], which motivated some later problems in noncommutative Borsuk-Ulam theory. Taghavi’s questions primarily concerned counting the number of odd self-adjoint elements required to produce an invertible square-sum. It should be noted that this is not the only interpretation of noncommutative Borsuk-Ulam theory; references [62, 3, 40, 42, 10, 11] consider an approach focused on (non-)existence of equivariant morphisms. The full formulation of equivariant dimension, including a much more general context, appeared in [20]. The local-triviality dimension of an action is defined in [20, Definition 3.1], and it is a noncommutative analogue of the Schwarz genus. Moreover, it is related to the Rokhlin dimension of [24, Definition 1.1], as explored in [20, §5.2]. We recall the definitions of the local-triviality dimension in Section 3.
In this manuscript, we will typically focus on actions of ${\mathbb{S}}^{1}$ or its finite cyclic subgroups, for which equivalent versions of the dimensions we study are found in [12, §3]. Consider, for example, a unital $C^{*}$-algebra $A$ on which ${\mathbb{Z}}/2$ acts. The action decomposes $A$ into eigenspaces for eigenvalues $\pm 1$; elements in the 1-1 eigenspace are called odd and elements in the $1$ eigenspace are called even. The local-triviality (LT for short) dimension of this action is
@@
where we note that the value $\infty$ is possible. With this notation, the Borsuk-Ulam Theorem is equivalent to the claim that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(C({\mathbb{S}}^{n}))=n$, where ${\mathbb{Z}}/2$ acts on $C({\mathbb{S}}^{n})$ via the antipodal action. Because $C({\mathbb{S}}^{n})$ is commutative, there are adjustments to this dimension that make no difference here, but vastly change the computation for noncommutative $C^{*}$-algebras. For example, one could relax the condition on the sum-square of the $a_{i}$ to be merely invertible instead of $1$. Alternatively, one could strengthen the requirements by insisting that the $a_{i}$ are normal commuting elements of $A$. Of particular note is that finiteness of the local-triviality dimension implies freeness [20, Theorem 3.8], though this is not always equivalent.
Here, we will study how the local-triviality dimension and its variants behave with respect to deformations and fields of $C^{*}$-algebras, with a focus on noncommutative spheres and tori. We are interested in circumstances where the local-triviality dimensions are continuous, or not continuous, with respect to a deformation parameter. By construction of examples, we find that the most one can generally hope for is semicontinuity of the dimension, as explored in Section 3. Further, we extend work of [12] by finding free actions whose weak local-triviality dimension is infinite. In Section 4, we explore the $\theta$-deformed spheres and tori in more detail, with a focus on the rational case. In particular, viewing a rational noncommutative torus as a space of sections of a finite-dimensional vector bundle yields information about the strong local-triviality dimension. These results strengthen the noncommutative Borsuk-Ulam theory of [40] when the domain sphere is commutative.
@@ 2 Preliminaries
For what follows, $\mathrm{Prim}(A)$ denotes the primitive spectrum [44, §4.1.2] of a $C^{*}$-algebra $A$: the space of primitive ideals, equipped with the Jacobson (or hull-kernel) topology [44, §4.1.4]. Further, $Z(A)$ denotes the center of $A$, and $M(A)$ denotes the multiplier algebra of $A$. Following [47, §3], say, we write
@@
The Natsume-Olsen spheres, which we will also call $\theta$-spheres for brevity, have $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\cong{\mathbb{Z}}$ [36, Theorem 4.3], with a specific generator that is inspired by the commutative case.
@@
Recall (e.g. [36, p.1094]) that for the classical 3-sphere, the group $K_{1}(C({\mathbb{S}}^{3}))$ is generated by the element
@@
More generally, [36, §5] shows that the cyclic group $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\cong{\mathbb{Z}}$ admits a generator of matrix dimension $2^{n-1}$, which we will denote $Z_{\theta}$, such that each entry is a *-monomial. For each of the ${\mathbb{Z}}/k$ rotation actions defined above, the eigenspace of each entry is such that there exist unitary matrices $A,B\in M_{2^{n-1}}({\mathbb{C}})$, independent of $\theta$, such that $R(Z_{\theta})=AZ_{\theta}B$ (where $R$ is applied entrywise) and $A^{k}=B^{k}=I$. The antipodal action ($k=2$) is particularly simple; in this case each entry of $Z_{\theta}$ is odd, so $R(Z_{\theta})=-Z_{\theta}$. See [40, §3] for a slightly modified construction that chooses the monomial coefficients continuously.
The Borsuk-Ulam Theorem extends to the Natsume-Olsen spheres by [40, Corollary 4.12]; if $\phi:C({\mathbb{S}}^{2n-1}_{\theta})\to C({\mathbb{S}}^{2n-1}_{\omega})$ is ${\mathbb{Z}}/k$-equivariant, then the $K_{1}$ degree $m$ of $\phi$ is congruent to $1$ modulo $k$. In particular, the induced map $K_{1}(C({\mathbb{S}}^{2n-1}_{\theta}))\to K_{1}(C({\mathbb{S}}^{2n-1}_{\omega}))$ is nontrivial. However, the local-triviality dimensions of these actions are not all identical. In particular, [10, Proposition 3.21] shows that the $\theta$-sphere $C({\mathbb{S}}^{2n-1}_{\theta})$ such that each generator anticommutes has local-triviality dimension $1$. Since the local-triviality dimension is integer-valued and the commutative sphere has dimension 2n12n-1, this immediately shows that the local-triviality dimension fails to be continuous in the parameter $\theta$.
The $\theta$-spheres are related to a similar family of $C^{*}$-algebras called quantum (or noncommutative) tori ([21, §12.2] or [49, §1]). If $\theta$ is an antisymmetric $n\times n$ real matrix, the $C^{*}$-algebra $A^{n}_{\theta}$, or $C(\mathbb{T}^{n}_{\theta})$, is defined by generators $u_{1},\ldots,u_{n}$ and relations
@@
Since the set of $2$-dimensional antisymmetric matrices is one-dimensional, we abuse notation somewhat when $n=2$, and simply refer to $\theta=\theta_{12}\in\mathbb{R}$. The noncommutative tori and noncommutative spheres are related by polar decomposition [36, Theorem 2.5], which we recall in Section 4. In particular, the action $R$ discussed on $\theta$-spheres factors through a related action on the quantum tori:
@@
All $\theta$-spheres and noncommutative tori are known to be Rieffel deformations in the sense of [51, §10.2]. While it is not clear if the $\theta$-spheres or noncommutative tori form a continuous field over the set of antisymmetric matrices, [51, Theorem 8.13] implies that the field is separately continuous in each entry. In particular, the distinction between those two cases does not matter for $3$-spheres or $2$-tori. As we will discuss continuity of fields in more detail, we recall some basic definitions.
Recall [26, Definition 1.5] that for a locally compact Hausdorff space $X$, a $C_{0}(X)$-algebra is a $C^{*}$-algebra $A$ equipped with a non-degenerate morphism $C_{0}(X)\to Z(M(A))$. For us, $X$ will typically be compact and $A$ will be unital, so there will be little need to worry about units. We will also revert to the notation $C(X)=C_{0}(X)$ when $X$ is compact.
@@
For a group $\mathbb{G}$} and a locally compact Hausdorff space $X$, a $\mathbb{G}$$C_{0}(X)$-algebra is a $C_{0}(X)$-algebra equipped with an action of $\mathbb{G}$} by automorphisms that fix the structure morphism $C_{0}(X)\to Z(M(A))$. $\blacklozenge$
$C(X)$-algebras are one way to formalize the notion of a field of $C^{*}$-algebras [48, Definition 1.1 and Propositions 1.2 and 1.3], while [38, Theorem 2.3] equates the notion to that of an upper semicontinuous $C^{*}$-algebra bundle, generalizing [19, Definition VII.8.2].
@@
A $C(X)$-algebra AA is called a continuous field if for each $a\in A$, the mapping
@@ 3 Local-Triviality Dimension and Fields
If $(C({\mathbb{G}}),\Delta)$ is a compact quantum group with a coaction on a unital $C^{*}$-algebra $A$, then [20, Definition 3.1] defines the weak and plain local-triviality dimensions as follows. Here, we let $t$ denote the identity function in $C_{0}((0,1])$. We also note that since $C({\mathbb{G}})$ denotes a compact quantum group, it need not be commutative as a $C^{*}$-algebra, despite the notation used.
@@
The local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{LT}}(A)$ is the infimum of the set of $n\in\mathbb{N}$ such that there exist $\mathbb{G}$}-equivariant *-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ satisfying $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)=1$. $\blacklozenge$
@@
The weak local-triviality dimension $\mathrm{dim}^{{\mathbb{G}}}_{\mathrm{WLT}}(A)$ is the infimum of the set of $n\in\mathbb{N}$ such that there exist $\mathbb{G}$}-equivariant *-homomorphisms $\rho_{0},\ldots,\rho_{n}:C_{0}((0,1])\otimes C({\mathbb{G}})\to A$ with $\sum\limits_{j=0}^{n}\rho_{j}(t\otimes 1)$ invertible. $\blacklozenge$
The strong variant of the local-triviality dimension, found in [20, Definition 3.20], is defined in reference to the noncommutative join operation of [14, Definition 4.1] or [3, Definition 1.4]. Here $E_{n}^{\Delta}{\mathbb{G}}$ denotes the join of $n+1$ copies of $\mathbb{G}$}.
@@
A coaction $\delta:A\to A\otimes_{\text{min}}C({\mathbb{G}})$ is called free (see [18] or [3, (1.10)]) if
@@
Multiple characterizations of freeness are also found in [4, Theorem 0.4]. Moreover, finiteness of any one of the local-triviality dimensions implies freeness of the coaction by [20, Theorem 3.8]. We will primarily be interested in the case that $A$ is noncommutative, but $\mathbb{G}$} is an ordinary group (in particular, a compact abelian group). In this case, there is an equivalent condition to freeness found in [50, Definition 1.6] or [45, Definition 7.1.4]. If $\mathbb{G}$} is compact abelian and $\alpha$ is an action of $\mathbb{G}$} on $A$, then for each character $\gamma\in\widehat{{\mathbb{G}}}$, there is an associated eigenspace (or spectral subspace)
@@
This “asterisk” shape is star-convex, with center 0, generated by the $k$th roots of unity.
@@
Since $\operatorname{Star}_{2}=[-1,1]$, the elements $a_{i}$ used to compute ${\mathbb{Z}}/2$-dimensions are self-adjoints. As such, dimensions of ${\mathbb{Z}}/2$-actions are typically much easier to compute than other cases. Also, as the set of invertible elements of $A$ is open, but $\{1\}$ is closed, the propositions suggest that weak and plain local-triviality dimensions may behave differently for fields of $C^{*}$-algebras. At the very least, for $\mathbb{G}$}-$C(X)$-algebras, we have the following immediate estimate on the global dimensions in terms of their respective fiber-wise counterparts.
@@
For a compact group $\mathbb{G}$} and a $\mathbb{G}$}-$C(X)$-algebra AA, we have
@@
The ${\mathbb{Z}}/2$-action, on the other hand, is “twisted” conformant to the topology of $X$: the action on the fiber $A_{x}\cong M_{2}$ for a line
@@
so the $1$-eigenspace of $v_{x}$ will coincide with neither of the two eigenspaces of $u_{x}$ (with respective eigenvalues $\pm 1$). Identifying
@@
said eigenspaces of $u_{x}$ are nothing but $x$ and its antipodal point $x\in{\mathbb{S}}^{2}$. The map
@@
is thus a self-map of the sphere that has no fixed points and sends no point to its antipode, and no such map exists by [59, Corollary 1.24]. $\blacklozenge$
@@
Equip ${\mathbb{P}}^{1}$ with its complex analytic structure or, equivalently for the purpose of manipulating holomorphic bundles, its complex projective algebraic structure (see [54, §12]). The line bundle
$X={\mathbb{P}}^{1}\ni x\xmapsto{\quad}\text{$1$-eigenspace of }u_{x}$
is the universal subbundle [16, §3.2.3] on ${\mathbb{P}}^{1}$ and hence what algebraic geometers denote by ${\mathcal{O}}(-1)$ [22, Definition preceding Proposition 5.12]. Its orthogonal complement
@@
in the trivial rank-2 bundle on ${\mathbb{P}}^{1}$ is then ${\mathcal{O}}(1)$, as the degrees [22, Exercise II.6.12] of the two bundles must add up to 0. A choice of non-zero self-adjoint odd $v_{x}$ for $x$ ranging over some open $U\subseteq{\mathbb{P}}^{1}$ means precisely an identification ${\mathcal{O}}(-1)\cong{\mathcal{O}}(1)$ over $U$, or a trivialization of
@@
Such trivializations exist over any proper open subset of ${\mathbb{P}}^{1}$ (e.g. the complement of a point), per the standard classification [39, Theorem 2.1.1] of line bundles on ${\mathbb{P}}^{1}$. It follows that there are trivializations
@@
A salient example of a $C(X)$-algebra is that of sections of an $n\times n$ matrix-algebra bundle $\mathcal{M}$} on $X$ in the sense of [25, Definition 18.1.3]: a vector bundle ${\mathcal{M}}\to X$ with local trivializations ${\mathcal{M}}|_{U}\cong U\times M_{n}$ respecting the $C^{*}$-algebra structures along the fibers. There will be no harm in allowing $n$ to vary, as $X$ simply decomposes into clopen subspaces over which $n$ is constant.
Having to work with such trivializations extensively, we need some language to express the extent to which they exist. To that end, the following notions specialize the (numerical) $\mathbb{G}$}-index ([63, Definition 3.13], [31, Definition 6.2.3]) of a $\mathbb{G}$}-action on a space $X$, which is typically compact. That invariant can be recast via [13, Theorem 3.4] (where the term is co-index instead) as the smallest $n$ for which $X/{\mathbb{G}}$ admits an open cover by $U_{0},\ldots,U_{n}$, above which the action induces trivial principal $\mathbb{G}$}-bundles.
@@
    Parts 1 and 2 of Definition 3.9 are indeed instances of numerical $\mathbb{G}$}-indices (at least along the clopen subsets of $X$ where the ranks of the vector/matrix bundles involved are constant), with $\mathbb{G}$} either a unitary or projective unitary group. As explained in [25, Assertion 18.3.4], rank-$n$ vector bundles and $n\times n$ matrix bundles “are” nothing but principal $U(n)$- and $PU(n)$-bundles, respectively.
@@
Because $\operatorname{\mathrm{qt}}$ is assumed finite, each ${\mathcal{E}}_{i,j}$ has nonzero vector space dimension at each $x\in X$. If $i$ is fixed and $x\in U_{i}$ is fixed, consider the smallest such eigenspace $W_{0}:=({\mathcal{E}}_{i,j_{0}})_{x}$. It is possible to define a partial isometry on $M_{n_{x}}$ that maps $W_{0}$ isometrically onto a subspace $W_{1}\subseteq({\mathcal{E}}_{i,j_{0}+1})_{x}$ (which is guaranteed to be orthogonal to $W_{0}$ because the eigenspaces are orthogonal), maps $W_{1}$ isometrically onto some $W_{2}\subseteq({\mathcal{E}}_{i,j_{0}+2})_{x}$, and so on, eventually returning to $W_{k}=W_{0}=({\mathcal{E}}_{i,j_{0}})_{x}$. Further, map every element of $\left(\bigoplus_{m=0}^{k-1}W_{m}\right)^{\perp}$ to 0. This procedure produces a partial isometry $\alpha$ in the generating spectral subspace of the action of ${\mathbb{Z}}/k$ on this fiber. In fact, $\alpha$ is normal, as the initial space and final space are both equal to the direct sum $\bigoplus_{m=0}^{k-1}W_{m}$, and it also follows from the construction that $\alpha^{k}$ is equal to the projection onto that same direct sum. The subspaces $W_{m}$ for $m>0$ may be proper subspaces of the corresponding eigenspaces, but because $\operatorname{\mathrm{qt}}$ is defined as the (largest) quotient spread, we need at most $\operatorname{\mathrm{qt}}$ such operators $\alpha_{1},\ldots,\alpha_{\operatorname{\mathrm{qt}}}$ to guarantee full coverage of each eigenspace. That is, $\sum\limits_{s=1}^{\operatorname{\mathrm{qt}}}\alpha_{s}^{*}\alpha_{s}$ is an invertible element of $M_{n_{x}}$.